### v1

# Deep Common Feature Mining for Efficient Video Semantic Segmentation

Abstract
Recent advancements in video semantic segmentation have made substantial progress by exploiting temporal correlations. Nevertheless, persistent challenges, including redundant computation and the reliability of the feature propagation process, underscore the need for further innovation. In response, we present Deep Common Feature Mining (DCFM), a novel approach strategically designed to address these challenges by leveraging the concept of feature sharing. DCFM explicitly decomposes features into two complementary components. The common representation extracted from a key-frame furnishes essential high-level information to neighboring non-key frames, allowing for direct re-utilization without feature propagation. Simultaneously, the independent feature, derived from each video frame, captures rapidly changing information, providing frame-specific clues crucial for segmentation. To achieve such decomposition, we employ a symmetric training strategy tailored for sparsely annotated data, empowering the backbone to learn a robust high-level representation enriched with common information. Additionally, we incorporate a self-supervised loss function to reinforce intra-class feature similarity and enhance temporal consistency. Experimental evaluations on the VSPW and Cityscapes datasets demonstrate the effectiveness of our method, showing a superior balance between accuracy and efficiency.
Figure 1: Trade-off illustration: various semantic segmentation methods using MiT backbones on VSPW validation set [28], including SegFormer (image baseline) [43], CFFM [36], MRCFA [37], and DCFM (Ours). By adjusting the keyframe interval $K$ during inference, DCFM achieves impressive speed (80 FPS @ $K$=10) while maintaining high level of accuracy (45.4% mIoU).
1 Introduction
Semantic segmentation is a prominent dense prediction task in the field of computer vision, involving the assignment of pixel-wise category labels to images. Significant progress in this field has been propelled by the exploration of robust deep neural backbones [32, 54, 14, 40, 43, 9], elaborate context aggregation strategies [5, 6, 7, 50, 47, 24, 43, 9], and additional boundary supervision [5, 33, 48]. In recent years, the proliferation of real-world applications, ranging from autonomous driving to robot sensing, has intensified the focus on video semantic segmentation (VSS). The success of this task extends beyond achieving high accuracy [13, 30, 36, 37, 2], necessitating a trade-off between inference efficiency [46, 12, 31, 55, 45, 15, 1] and temporal consistency [25, 49].
While it is straightforward to extend the capabilities of contemporary image models to VSS, the naive adaptation to every video frame introduces considerable computational overhead and neglects the inherent temporal correlations in videos, which are crucial for understanding scenes with lower uncertainty and enhancing temporal smoothness [25, 49]. To address this, methods like LMANet [30], CFFM [36], and MRCFA [37] explicitly aggregate information over multiple frames, utilizing Transformers [38] or spatial-temporal attention for context encoding. Despite their effectiveness, the heavy computational cost and complex temporal modeling hinder practical deployment [13, 39, 19, 30, 36, 37]. As illustrated in Figure 1, achieving high accuracy with CFFM and MRCFA in certain experimental environment leads to a slow inference speed of approximately 6 FPS.
On the flip side, the temporal continuity in videos presents an opportunity to enhance efficiency. Acknowledging the frequent redundancy in content across adjacent frames, a promising approach is to reuse high-level features extracted from selected keyframes, striking a refined balance between inference speed and accuracy [31, 55, 34, 45, 1, 16]. This can be achieved by leveraging a pre-trained optical flow model for feature warping [17] or a well-designed propagation module [31, 1, 16] to transmit features or segmentation maps from keyframes to adjacent ones, significantly reducing redundant computation. However, existing keyframe-based methods tend to excessively rely on the propagation process, leaving accurate feature prediction unsolved, especially in complex scenarios involving large motions and severe occlusions [31, 52, 49]. Furthermore, propagating unstable features tends to accumulate errors [44].
In this study, our emphasis is on achieving efficient VSS by learning a discriminative representation to assist adjacent frames in perceiving the scene without resorting to time-consuming re-calibration operations like optical-flow-based warping. To address both aspects, we propose decomposing the entangled feature into two complementary components: (1) the common representation, providing essential information for segmentation and directly reusable across frames without feature propagation. This corresponds primarily to high-level semantics that remain relatively constant over time; and (2) the independent representation, easily extractable to offer frame-specific clues for accurate segmentation. Figure LABEL:fig:moti visualizes the distinctions between our approach and the existing ones.
To achieve this goal, we propose a novel approach called Deep Common Feature Mining (DCFM). In DCFM, we organize the deep and shallow layers of the backbone network into stages to capture common and independent representations, respectively. A lightweight fusion module is designed to integrate spatially misaligned multi-level features, ensuring a comprehensive representation and high efficiency. However, a notable challenge in common feature mining arises from sparsely annotated training data, as exemplified by datasets like Cityscapes [10], where only one frame is labeled in a video. To address this limitation and facilitate the learning of a high-level feature containing sufficient common information across consecutive frames, we propose a symmetric training strategy, alternately considering labeled frames as key-frames and non-keyframes. Additionally, a self-supervised loss reinforces intra-class feature similarity, considering both static and dynamic contents. Crucially, our model operates without re-calibration, contributing significantly to improved result robustness and achieving an extremely fast speed on non-key frames.
Our main contributions include: (1) a novel framework for efficient VSS featuring a common feature mining architecture and a symmetric training strategy; (2) a novel self-supervised consistency loss to enhance representation robustness, accommodating both static and rapidly evolving contents; and (3) substantial speed-ups over baseline models and an improved speed-accuracy trade-off compared with the current state-of-the-art.
2 Related work
2.1 Image semantic segmentation
Fully Convolutional Networks (FCN) [27] paved the way for the application of Convolutional Neural Networks (CNNs) in image semantic segmentation. Recognizing the necessity for informative spatial contexts, subsequent methods focused on refining network architectures to aggregate rich contextual information. Among them, the encoding-decoding structures have proven to be particularly effective [32, 54, 8, 3]. PSPNet [50], along with studies like DeepLabV3 [7] and OCRNet [47], introduced post-processing modules for enhanced spatial awareness. HRNet-V2 [40] utilized exchange blocks for multi-scale information fusion. The application of vision transformers in image tasks, as seen in SETR [53] and Segformer [43], further advanced this field. Additionally, MaskFormer and Mask2Former improved DETR [4] for segmentation tasks, steering away from pixel-by-pixel classification. Some methods also incorporated boundary supervision [5, 33, 48] to enhance accuracy.
Another notable trend aimed at designing efficient networks for practical deployment [46, 21, 22, 51]. In particular, STDC [12] removed structural redundancy from BiSeNet [46] and introduced a new backbone, considering features extracted by shallow layers as the output of the spatial path, thereby further reducing model complexity.
2.2 Video semantic segmentation
To exploit temporal correlations in videos, various methods [13, 18, 19, 29, 15, 39, 30, 36, 37, 2, 23] explicitly aggregate high-level features across frames for improved accuracy. Attention mechanisms [38] and vision transformers [11], recognized for their proficiency in long-term dependency modeling, are prominent in VSS. TDNet [15] and TMANet [39], for instance, use attention-based approaches for feature aggregation, while CFFM [36] employs a transformer to learn a unified representation of static and motion contexts. Despite their accuracy, these methods often demand substantial computational resources.
Another category of methods explores the advantages of video continuity, prioritizing a favorable speed-accuracy trade-off [31, 55, 34, 45, 41, 16]. Typically, a robust baseline segmentation model is selectively applied to keyframes, and the extracted high-level features are reused by subsequent non-key frames, effectively reducing average latency. However, an additional propagation module is generally required, introducing extra computational overhead.
The literature emphasizes temporal consistency, often achieved by aggregating multi-frame information during inference [29, 20, 35]. Additionally, [25] introduced a knowledge distillation-based method for consistent image semantic segmentation results, leveraging temporal loss. To address warping errors from optical flow, [49] proposed the Perceptual Consistency (PC) measure, providing a more accurate evaluation of temporal consistency. Notably, PC served as a regularizer during training to enhance temporal consistency. The VSPW dataset [28] contributes by providing frame-by-frame annotations, facilitating the calculation of video consistency (VC) based on ground-truth.
3 Method
This section describes the DCFM approach, delving into its network architecture, efficient inference pipeline, and the innovative training strategy with a specialized loss function.
3.1 Overview
We employ a contemporary image segmentation model as the backbone. As observed in Clock [34] and STDC [12], features of the deep layers of the backbone tend to capture high-level semantics, evolving gradually over time. In contrast, shallow features excel at depicting details with quicker update rates. During training, we emphasize this property through common feature mining, ensuring deep representations from keyframes can be effortlessly reused across frames during inference, eliminating the need for re-calibration. This achieves an optimal balance between inference speed and accuracy. To accomplish this, we group the deep and shallow layers of the backbone network into two stages to capture the common and independent representations, respectively. The symmetric training strategy alternates between a labeled keyframe and a non-key frame, facilitating high-level common feature mining (detailed in Section 3.2). Additionally, a proposed self-supervised loss minimizes perceptual distance for intra-class pixels, enhancing representation robustness (detailed in Section 3.3).
In Figure LABEL:fig:framework, we present the complete network architecture when the model is applied to a video. For a video comprising frames $\{x_{0},x_{1},...,x_{i},...\}$, where $x_{i}\in\mathbb{R}^{C\times H\times W}$, we select one keyframe for every $K$ frames. The shallow and deep layers of the backbone are organized to extract features from this keyframe $x_{i}$, expressed as:
$F^{i}_{co}=$ $~{}\text{Conv}(\text{Enc}(x_{i})),$ (1)
$F^{i}_{in}=$ $~{}\text{Enc}_{lo}(x_{i}),$ (2)
[…]
$F^{i}=$ FFM(Norm($~{}\text{FFM}(\text{Norm}(F^{i}_{co}),\text{Norm}(F^{i}_{in})),$ (3)
$\hat{y}^{i}_{I}=$ $~{}\text{Dec}(F^{i}),$ (4)
where $F^{i}\in\mathbb{R}^{C_{F}\times H_{F}\times W_{F}}$ denotes the fused representation of the keyframe $x_{i}$, Norm signifies the normalization operation, and Dec represents the segmentation head. Similar to many previous image segmentation studies [46, 12], a combination of low-level details and high-level semantics is utilized to enhance prediction accuracy.
Given sparse annotations in existing video segmentation datasets, we adopt a symmetric training strategy to reinforce the temporal coherence of the deep layers while ensuring that the shallow features contain rich frame-specific dynamics. At the inference stage, the independent feature $F^{j}_{in}$ of a non-key frame $x_{j}$ can be trivially extracted by a single feed-forward through the shallow layers, and the common features extracted from its two nearest keyframes, $x_{i}$ and $x_{k}$, are explicitly reused. The segmentation result $\hat{y}^{j}_{B}$ is efficiently obtained by:
$F^{i\rightarrow j}$ $=~{}\text{FFM}(\text{Norm}(F^{i}_{co}),\text{Norm}(F^{j}_{in})),$ (5)
$\hat{y}^{i\rightarrow j}$ $=~{}\text{Dec}(F^{i\rightarrow j}),$ (6)
$\hat{y}^{j}_{B}$ $=~{}\frac{1}{2}[\hat{y}^{i\rightarrow j}+\hat{y}^{k\rightarrow j}],$ (7)
where $F^{i\rightarrow j}$ denotes the feature obtained by fusing $F^{j}_{in}$ with $F^{i}_{co}$, and $\hat{y}^{i\rightarrow j}$ indicates its corresponding result. Similarly, $F^{k\rightarrow j}$ and $\hat{y}^{k\rightarrow j}$ can be computed. This approach leverages both the previous and subsequent keyframes to achieve more robust segmentation results.
Feature fusion module. The structure of FFM is illustrated in Figure LABEL:fig:ffm. Initially, we align the sizes of common and independent feature maps through interpolation. To alleviate computational burdens and prevent short-distance back-propagation from affecting training, we concatenate $F^{i}_{co}$ with half of the independent feature $F^{i}_{in}$ (or $F^{j}_{in}$). Then we employ only one convolution layer for feature fusion, ensuring minimal computational load.
3.2 Deep common feature mining
To extract common features between frames, we implement a symmetric training scheme guiding the backbone to learn an effective high-level representation with abundant common segmentation clues from adjacent frames. As illustrated in Figure 2, each training iteration involves two neighboring frames, $x_{l}$ and $x_{u}$, where $x_{l}$ is labeled, and $u\in\{l-1,l+1\}$. Initially, the labeled frame $x_{l}$ serves as the keyframe, and $\hat{y}^{l}_{I}$ can be obtained by formulas (1) - (4). The basic cross-entropy loss is applied, formulated as:
$L_{I}=\text{CE}(\hat{y}^{l}_{I},y^{l}),$ (8)
where $y^{l}$ represents the corresponding ground-truth and CE indicates the loss function.
The primary challenge arises from determining how to supervise the feature of the unlabeled frame. To enhance the coherence of the deep features between these two frames, we introduce implicit supervision on the unlabeled frame $F^{u}_{co}$ by reusing it. In particular, the labeled frame $x_{l}$ is repurposed as a non-key frame, while the unlabeled frame $x_{u}$ functions as its closest keyframe, providing the deep common representation. This allows us to obtain $\hat{y}^{u\rightarrow l}$ through the formulas (1)-(2) and (5)-(6). As such, the supervision applied to the non-key frame $x_{l}$ can be expressed as:
$L_{B}=\text{CE}(\hat{y}^{u\rightarrow l},y^{l}).$ (9)
Minimizing the loss term (9) strengthens the consistency of common representations between neighboring frames. The alternate treatment of labeled frames as keyframes and non-key frames achieves deep common feature mining on sparsely annotated data. To facilitate learning, we further introduce intermediate supervision on the fused features, as detailed in the subsequent section.
Figure 2: Illustration of the common feature mining process. A labeled frame $x_{l}$ serves alternately as both the keyframe and non-key frame. This cyclic treatment facilitates implicit supervision of the deep features in an unlabeled neighboring frame $x_{u}$.
3.3 Self-supervised consistency loss
To alleviate the training complexity, we adopt a two-frame per-iteration training strategy. However, this choice restricts the explicit use of long-term supervisory signals to enhance the temporal consistency of segmentation results, as employed in [25, 49]. To address this limitation, we introduce a straightforward yet potent self-supervised temporal consistency loss, which can be seamlessly integrated into our framework. Specifically, this loss focuses solely on positions where predicted category labels remain constant over time, aiming to refine intra-class feature similarity. The formulation is as follows:
$M_{inter}=$ $~{}\mathbf{1}_{\{0\}}[\text{argmax}(\hat{y}^{u}_{I})-\text{argmax}(\hat{y}^{l}_{I})],$ (10)
$L_{C}=$ $~{}\text{MSE}(M_{inter}\cdot F^{u},M_{inter}\cdot F^{l}).$ (11)
where $M_{inter}\in\{0,1\}^{H_{F}\times W_{F}}$ denotes the intersection of two predicted semantic segmentation maps, $\hat{y}^{u}_{I}$ and $\hat{y}^{l}_{I}$. The term $MSE$ represents Mean Square Error. Minimizing this equation effectively reduces the feature distance between the corresponding positions of the two frames, for the locations where the predicted category labels remain consistent. Importantly, we refrain from pushing features apart at positions where the predicted results change, recognizing that features at these locations are less stable.
Figure 3: Illustration of the proposed consistency loss. We minimize the distance between $F^{u}_{p}$ and $F^{l}_{p}$, where $p\in M_{inter}$. On a moving object, this operation is akin to minimizing the distance between $F^{u}_{p1}$ and $F^{u}_{p3}$.
Figure 3 visually depicts the self-supervised consistency loss $L_{C}$, which comprehensively considers both static and dynamic contents. For static objects, $L_{C}$ guides the model to learn a more stable feature representation over time, enhancing temporal consistency. For moving objects, it encourages the model to learn a more similar feature representation for pixels at different positions within the same category. Taking the car in Figure 3 as an example, the pixel at $p_{1}$ in Frame $x_{l}$ moves to $p_{3}$ in Frame $x_{u}$, and their features naturally exhibit similarity. By narrowing the feature distance between $F^{u}_{p1}$ and $F^{l}_{p1}$ at position $p_{1}$, the distance between $F^{u}_{p1}$ and $F^{u}_{p3}$ is reduced, reinforcing intra-class similarity. Overall, the self-supervised consistency loss $L_{C}$ empowers the model to acquire more robust feature representations with heightened perceptual consistency.
In summary, the joint loss is expressed as:
$L=L_{I}+\lambda_{B}L_{B}+\lambda_{C}L_{C}.$ (12)
where $\lambda_{B}$ and $\lambda_{C}$ are the trade-off parameters.
3.4 Adaptive keyframe scheduling
Previous research [31, 45, 55] has made valuable contributions to adaptive keyframe scheduling (AKS), allowing keyframe-based VSS methods to adapt to dynamic scenes with varying motion levels instead of using a fixed keyframe scheduling. While DCFM does not explicitly integrate a keyframe scheduling module, we use a simple AKS scheme [34] to enhance flexibility. We calculate a score between each frame and the previous keyframe during inference, defined as $s_{j}=\text{mean}(|x_{i}-x_{j}|)$, where $x_{i}$ is the previous keyframe, and $j>i$. We traverse through $x_{j}$ until the score $s_{j}$ exceeds a predefined threshold $S$ or reaches the end of the video. This adaptive keyframe selection strategy allows DCFM to adapt to varying motion levels, providing flexibility in handling dynamic scenes.
Methods Backbone mIoU $\uparrow$ wIoU $\uparrow$ $\text{mVC}_{8}$uparrow $\text{mVC}_{16}$uparrow Params $\downarrow$ GFLOPS $\downarrow$ FPS $\uparrow$
DeepLabv3+ [7] ResNet-101 34.7 58.8 83.2 78.2 62.7M 379.0 9.25
UperNet [42] ResNet-101 36.5 58.6 82.6 76.1 83.2M 403.6 16.05
PSPNet [50] ResNet-101 36.5 58.1 84.2 79.6 70.5M 401.8 13.84
OCRNet [47] ResNet-101 36.7 59.2 84.0 79.0 58.1M 361.7 14.39
MPVSS [41] Swin-T 39.9 62.0 85.9 80.4 114.0M 39.7 32.86
MPVSS [41] Swin-S 40.4 62.0 86.0 80.7 108.0M 47.3 30.61
Mask2Former [9] Swin-T 41.2 62.6 84.5 80.0 47.4M 114.4 17.13
Mask2Former [9] Swin-S 42.1 63.1 84.7 79.3 68.9M 152.2 14.52
SegFormer [43] MiT-B0 32.9 56.8 82.7 77.3 3.8M 11.4 76.2
SegFormer [43] MiT-B1 36.5 58.8 84.7 79.9 13.8M 21.6 67.3
SegFormer [43] MiT-B2 43.9 63.7 86.0 81.2 24.8M 34.2 44.1
SegFormer [43] MiT-B5 48.2 65.1 87.8 83.7 82.1M 102.3 13.9
CFFM [36] MiT-B2 44.9 64.9 89.8 85.8 26.5M 149.6 11.4
MRCFA [37] MiT-B2 45.3 64.7 90.3 86.2 27.3M 142.0 12.3
DCFM (Ours, $K$=2) MiT-B2 43.7 63.7 87.7 83.2 24.8M 22.9 67.6
DCFM (Ours, $K$=6) MiT-B2 42.4 62.6 90.4 86.2 24.8M 12.3 113.6
CFFM [36] MiT-B5 49.3 65.8 90.8 87.1 85.5M 433.5 4.9
MRCFA [37] MiT-B5 49.9 66.0 90.9 87.4 84.5M 414.4 5.3
DCFM (Ours, $K$=2) MiT-B5 48.2 65.5 89.0 85.0 82.1M 57.0 25.4
DCFM (Ours, $K$=6) MiT-B5 46.5 64.1 91.7 88.0 82.1M 23.6 58.8
Table 1: Comparison of accuracy (mIoU and wIoU), temporal consistency (mVC), parameter number, GFLOPs, and inference speed (FPS) on the VSPW [28] validation dataset. Figure LABEL:fig:sota-vs provides a visual illustration, focusing on two dimensions: mIoU and FPS. $K$ indicates the keyframe interval during inference.
4 Experiments
4.1 Datasets and evaluation metrics
VSPW. The VSPW dataset [28] comprises a total of 2,806 videos, distributed across training (2,806 videos), validation (343 videos), and test sets (387 videos). The dataset encompasses 124 semantic categories and has a frame rate of 15 FPS. The frame resolution is $480\times 853$ pixels, providing detailed frame-by-frame annotations.
[…]
Evaluation metrics. We evaluate our approach using mean Intersection over Union (mIoU) and frequency-weighted Intersection over Union (wIoU) metrics on annotated frames. Additionally, we measure inference speed in frames per second (FPS). For assessing temporal consistency, we use mean Video Consistency (mVC) [28].
4.2 Implementation details
During training, data augmentation techniques such as random rotation, random resizing, random cropping, color jitter, and random horizontal flipping are implemented. The crop size is set to $480\times 480$ for VSPW and $768\times 1536$ pixels for Cityscapes. At the inference stage, the entire frame is fed into the model on a single Tesla V100 GPU, avoiding the sliding window input. Frame sizes are resized to $480\times 853$ for VSPW and $512\times 1024$ for Cityscapes when using MiT as the backbone. The default interval $K$ between keyframes is set to 2. Additional experimental details are available in the supplementary materials.
Methods Backbone mIoU $\uparrow$ FPS $\uparrow$
DFF [55] ResNet-101 69.2 5.6
GRFP [29] ResNet-101 69.4 3.2
DVSN [45] ResNet-101 70.2 11.5
ETC [25] ResNet-18 71.1 9.5
MGC [23] ResNet-18 73.7 20.6
Accel [18] ResNet-50 74.2 1.5
DCNet [1] ResNet-101 75.6 6.3
TDNet [15] ResNet-18 76.4 11.8
FCN [27] ResNet-101 76.6 2.8
LVS [31] ResNet-101 76.8 5.8
PSPNet [50] ResNet-50 77.0 7.8
LMA [30] ResNet-101 78.5 1.3
DCFM (Ours) ResNet-18 74.0 30.1
DCFM (Ours) ResNet-50 77.2 13.2
DCFM (Ours) ResNet-101 78.0 5.6
SegFormer [43] MiT-B0 71.9 58.5
MRCFA [37] MiT-B0 72.8 33.3
CFFM [36] MiT-B0 74.0 34.2
SegFormer [43] MiT-B1 74.1 46.8
MRCFA [37] MiT-B1 75.1 21.5
CFFM [36] MiT-B1 75.1 23.6
DCFM (Ours) MiT-B1 73.1 72.2
DCFM* (Ours) MiT-B1 74.4 47.2
Table 2: Performance comparison in terms of accuracy (mIoU) and inference speed (FPS) on the Cityscapes [10] validation dataset. Methods, excluding ours, are arranged based on mIoU. DCFM* indicates a variant of our model with a different input resolution ($640\times 1280$). Results are visualized in Figure LABEL:fig:sota-cs.
4.3 Comparison with previous approaches
We comprehensively compare our method with the existing ones on VSPW [28]. Results in Table 1 highlight the notable performance of our models, particularly DCFM-B2 and DCFM-B5. Compared to the image baseline (SegFormer [43]), DCFM-B2 and DCFM-B5 improve FPS from 44.1 and 13.9 to 67.6 and 25.4 (speed up 1.5 times and 1.8 times), respectively, while maintaining almost the same accuracy with the keyframe interval $K$ = 2. In comparison to MRCFA-B5, the current state-of-the-art method on VSPW in terms of accuracy, DCFM-B5 ($K$ = 2) exhibits a slight decrease of -1.7% in mIoU, but it achieves around fivefold increase in FPS (from 5.3 to 25.4). By adjusting $K$ to 6 during inference, our DCFM-B5 model achieves a significantly faster speed. Despite the slight decrease in mIoU, it still outperforms MRCFA-B2 and CFFM-B2 in terms of both accuracy and inference speed, despite having a larger parameter count. The results visualized in Figure LABEL:fig:sota-vs further demonstrate superior temporal consistency (mVC) and a more favorable trade-off between accuracy and efficiency achieved by the proposed approach.
We extend our model comparison to the Cityscapes [10], detailed in Table 2 and Figure LABEL:fig:sota-cs. DCFM-50 outperforms contemporary feature aggregation methods like TDNet, which prioritizes high accuracy while utilizing multiple distributed networks for enhanced inference speed simultaneously. Notably, our method achieves a superior mIoU while maintaining a higher inference speed. In comparison to the method with the highest mIoU score, LMA [30], our DCFM-101 exhibits a 0.5% lower mIoU but is 4.3 times faster. When compared with other keyframe-based techniques aimed at accelerating segmentation models, we surpass LVS [31] and DCNet [1] in both accuracy (77.2% mIoU) and speed (13.2 FPS) with a smaller backbone (ResNet-50 [14]). These two methods also leverage different layers of the backbone network for feature reuse like ours, while our approach stands out by decoupling keyframe information into the common and independent parts. Among the methods using MiT [43] as the backbone, DCFM is both faster and more accurate than the SegFormer [43] baseline. Compared with CFFM [36] and MRCFA [36], the most accurate methods on VSS, DCFM achieves 0.7% less mIoU but about twice as fast.
Results on both datasets affirm the effectiveness of DCFM. Further insights into the advantage of such feature decoupling are detailed in Section 4.4.
4.4 Ablation study
Effectiveness of feature decomposing. By decomposing features into common and independent representations, our approach distinguishes itself from other keyframe-based methods. In Table 3, we perform ablation study by omitting one of the feature components in keyframes or non-key frames during both training and testing. As shown in the first row, excluding the shallow features $F^{key}_{in}$ of the keyframe causes our framework to revert to the structure of previous feature propagation-based methods, as depicted in Figure LABEL:fig:moti_b. The observed significant decrease in accuracy suggests a misalignment between the keyframes and the corresponding non-keyframes. However, instead of applying a re-calibration operation on the utilized feature to address this issue, our framework mitigates misalignments by mining common information across frames. Visualization of segmentation results are available in the supplementary material. Comparing the results of the ablation model in the second row with those obtained by the full model, it is evident that the learned common feature is highly informative, containing most of the necessary information for the non-key frames. However, utilizing the common feature $F^{key}_{co}$ provided by the keyframe alone on non-key frames is suboptimal. This is primarily because it neglects the quickly changing frame-specific dynamics, and such complementary information is well captured by the independent feature in our full model.
Key Non-key $\text{mIoU}_{k}$ $\text{mIoU}_{n}$ $\text{mIoU}_{K=2}$
$F^{key}_{in}$ $F^{key}_{co}$ $F^{non}_{in}$ $F^{key}_{co}$
42.0 41.8 41.9
41.8 40.8 41.3
43.9 43.6 43.7
[…]
$L_{C}$ mIoU $\text{mVC}_{8}$ $\text{mVC}_{16}$
w/o 42.3 87.3 82.5
w/ 43.7 87.7 83.2
Table 4: Quantitative ablation study on the consistency loss function, achieved on the VSPW [28] validation dataset.
Effectiveness of consistency loss. Table 4 illustrates that the inclusion of $L_{C}$ significantly enhances segmentation accuracy and temporal consistency. Figure 4 visually evaluates the impact by calculating cosine similarity between each pixel and its surrounding eight pixels in the feature space (output of FFM), indicating that the learned features are less susceptible to rapidly changing details in areas with the same semantic category when applying $L_{C}$.
Modules Methods SegFormer-B5 DCFM-B5 (Ours)
Key Non-key
Backbone $\text{Enc}_{lo}$ 2.5ms 2.5ms 2.5ms
$\text{Enc}_{hi}$ 68.6ms 68.6ms -
Feature Aggregation - 1.1ms 1.6ms
Segmentation Head 0.7ms 0.7ms 1.7ms
Total 71.8ms 72.9ms 5.8ms
Table 5: Comparison of the inference speed on VSPW [28].
Figure 4: Visualization of the cosine similarity between each pixel and its surrounding ones. Patches in yellow and red boxes show the effects of $L_{C}$ on intra- and inter-class pixels, respectively.
Figure 5: Results of adaptive keyframe scheduling on VSPW [28].
Analysis on inference speed. As shown in Table 5, for a keyframe, DCFM only uses an additional lightweight FFM to integrate features before the header, the inference time increases by only 1.1ms compared with the SegFormer-B5 [43] baseline. Regarding a non-key frame, the computing capacity of the shallow network has only 1.48 GFLOPS, thus DCFM achieves a very fast speed of 5.8ms (172 FPS) by reusing the keyframe common feature, constituting only 8% of the keyframe inference time.
[…]
5 Conclusion
This study addresses challenges in efficient video semantic segmentation. The proposed DCFM method aims to enhance accuracy and temporal consistency while overcoming computational bottlenecks and feature propagation in existing approaches. It decomposes the intricate features into two components, i.e., common and independent representations, where the former can be directly reutilized by neighboring frames without additional re-calibration, contributing to an optimized trade-off between accuracy and speed. A symmetric training strategy and a self-supervised consistency loss are introduced, enhancing the ability to handle sparsely annotated data, offering potential applications in other video analysis tasks.